# Reranking Social Media Feeds: A Practical Guide for Field Experiments

### v1

@@ 3.3. Content Editing
In addition to re-ranking the content of the feed, researchers can operate directly on individual posts. This intervention can involve direct manipulation of the post, such as editing the posts’ visible social metrics (e.g., likes, comments, and shares), the appearance of the post (e.g., making some aspects of the post prominent), changing the attachments (e.g., replacing a link or a picture), or modifying the text. These interventions may be important in addressing research questions on how the platform’s design or content can affect the users. For example, the tone of existing posts can be subtly reframed to reflect a more positive sentiment or mitigate aggressive language. This intervention can allow researchers to explore how language choices in social media posts may influence user behavior, including interactions, affective experiences, discourse patterns, misinformation, or hate speech. Depending on the complexity of the desired edits, the modification can be achieved with simple dictionary substitution or complete reframing, e.g., leveraging LLMs to generate high-quality and contextual text.
Summary. This section summarises the opportunities and challenges associated with down-ranking, up-ranking, and editing feed content. Down-ranking is generally the simplest intervention as it requires operating on content already available in the feed, whereas up-ranking, in the absence of access to the complete inventory, may require a more careful design to find or generate the posts to up-rank. Content editing can be combined with the previous methods and represents an effective strategy for testing the impact of social metrics, language, or post appearance. Furthermore, thanks to browser extensions’ capabilities, these modifications can be combined with additional interventions. Some examples include extending the platform’s functionality with widgets like the Twitter/X Community Notes labels or adding visual indications on some post characteristics, such as a warning about language toxicity. These interventions can be combined to achieve more complex experimental designs that fit particular research goals.
Finally, all these interventions require researchers to pay attention to the distribution of posts in the feed of the target population, potentially with some repercussions for the recruitment strategy. For example, manipulating a single post may produce minimal exposure to the treatment condition and the effect of the intervention may be too small to detect (Guess, 2021). This problem is particularly impactful in situations when users interact with only a small fraction of the content (Wang et al., 2016), typically at the top of the feed (Bakshy et al., 2015) or when the types of posts that the researchers intend to edit are rare.
@@ 4. Measurement
In this section, we describe a set of approaches to measuring the feed intervention’s impact. We summarise three alternatives that enable different insights into the effect of the feed manipulation: momentary measures, longitudinal surveys, and engagement signals.
@@ 4.1. Ecological Momentary Assessments
When researchers are interested in measuring the momentary impact of a feed intervention, a suitable design is to add surveys directly to the feed. These in-feed surveys, inspired by the approach developed in clinical psychology called Ecological Momentary Assessment (Shiffman et al., 2008), allow researchers to capture real-time reactions directly in the context where the intervention is applied (Backstrom, 2016). In the case of social media, this strategy can elicit direct feedback on the intervention at the right time when users scroll through the social media feed. Researchers can ask highly contextualized questions such as “Are you interested in this post?” or “How does this post make you feel?”. This method can be used to measure a momentary impact of the intervention that may be too short to be measured with long-term surveys, or that is hard to measure when the question is posed without context. The insertion of these surveys can be triggered by specific events, such as when a specific post is displayed, added at regular intervals, or shown at random times while the user consumes their feed. The appearance of the survey widget may vary, from attention-grabbing designs such as popups or modal windows that can interrupt the scrolling, to more subtle designs integrated as a special post in the regular feed. If integrated with the regular feed, one practical challenge of in-feed surveys is making them visible and clear enough for participants to know these questions are part of the study. If researchers require that the insertions get noticed, we recommend using colors and animations that stand out from the default design. The design must also be adapted to the different templates (i.e., dark mode theme) that the social media platform supports. Finally, when deciding where to place the in-feed survey, researchers must carefully consider the goals of their measurement. Participants answering the survey will be exposed to the content that precedes it in the feed, and might be influenced by other content visible in their viewport.
@@ 4.2. Survey methodology
Researchers interested in measuring the cumulative effects of the treatment can rely on a standard pre-post survey design. In such a setup, participants are presented with the same survey before (pre-) and after (post-) the experiment. Examples of attitudinal shifts that can be caused by feed interventions include changes in affective polarization, often measured using the feeling thermometer, or opinion change, typically measured with Likert-scale surveys. These surveys are assigned at the start and the end of a study and require a seamless integration into the experimental design. Pre-surveys intended to be completed at the beginning of the experiment can also be used to assign participants to specific experimental conditions or decide whether they qualify to be enrolled in the experiment. Attention checks and validation questions can be valuable tools for identifying early participants who may be less attentive and contribute unreliable data.
A similar approach, reassembling a diary study technique (Bolger et al., 2003), consists of regular assessments throughout the experiment. This method relies on administering surveys at regular intervals, e.g., daily or weekly. This design may be recommended for different reasons, such as measuring the progressive change in the outcomes of interest or because of a study design requiring participants to be exposed to different experimental conditions (e.g., crossover or stepped-wedge design).
Implementing a strategy to administer surveys at regular intervals can present additional challenges, especially when the users are not online on the social media platform. Researchers need to establish a notification system through the browser extension, implement regular messaging on the recruitment platform, or may need to distribute a survey link by establishing extra communication channels like email or text messaging. This last approach requires participants’ consent to collect potentially identifiable information, which we discuss in more detail in Sec. 5.4. Another challenge of this setup is ensuring that participants consume enough content in time between surveys to meaningfully measure the impact of the intervention. One potential approach, valid in the case of repeated assessments, is to invite the participants to use social media and deliver the survey only after they spent enough time on the platform or have been exposed to a predefined number of posts.
@@ 4.3. Engagement signals
Implementing a feed experiment using a browser extension allows researchers to go beyond self-reported measures and make very granular behavioral measurements that can provide further insights into the effects of the intervention. This includes metrics such as time spent, engagement (e.g., clicks, likes/favorites, reports), navigation patterns, conversion rates, and other metrics (Cunningham et al., 2024) that are imperative to the business objectives of the platforms. Access to such measures also allows researchers to evaluate the trade-offs the proposed intervention introduces. For instance, removing some content entirely (e.g., all political posts) may have an impact on an outcome of interest (e.g., affective polarization) but may disrupt the engagement in a way that makes the adaption not attractive for social media companies. Reporting these metrics may make social media platforms more receptive to integrating the findings into their products.
@@ 5. Implementing a Browser Extension
In this section, we focus on the implementation details of running social media feed experiments with a browser extension and summarize various technical challenges. We offer practical recommendations and release source code  [^2] https://github.com/StanfordHCI/FeedMonitor—in Javascript—that can be used as a blueprint to implement feed experiments on X. Our examples are based on Google Chrome but can easily be adapted to other browsers. The rest of this section assumes familiarity with key concepts of web architecture and extension development.
@@ 5.1. Feed interventions
In this section, we present some key details on the implementation of the core functionality of the browser extension. Although these extensions can change many aspects of the experience on a website, here we focus specifically on some edits they can support on social media feeds. Browser extensions are typically written in languages supported by the browser, such as Javascript, HTML, and CSS, and they can employ different methods to apply the edits. A common strategy involves modifying the rendered page by manipulating its Document Object Model (DOM). This method has the advantage of simplifying the identification of the element involved in the desired edit, and it gives the extension access to the content after the frontend scripts have already processed it. However, despite this approach being the most straightforward, for some scenarios, such as augmenting the feed, intervening at the network level may be more effective.
@@
In this case, we recommend adapting XMLHttpRequest to support the customization of the server requests. This technique offers the advantage of intercepting and modifying the original incoming data before its rendering on the page, minimizing the need for direct interaction with the actual page interface. Accessing the communication with the server allows logging many client-side actions shared with the server (e.g., like and share) without adding a listener for every event. The downside is the need to gain a deeper understanding of the communication protocol between the browser and the server. Finally, alternative approaches, such as manipulating the in-memory data structure used to render the content, are possible, but depending on the framework used by the front end (i.e., React), it may introduce implementation-specific complications.
@@
Implementation details. Pragmatically, to customize the behavior of the XMLHttpRequest native object, we need to run the overriding code as soon as the main page is loaded. Since behavior customization must happen within the scope of the social media page and not in the content script, we need to inject the modification script directly into the main page. This requirement demands a slight change from the common pattern used to develop browser extensions, as the script must be loaded in the Manifest file as a resource and not as a content script. Then, the content script can inject the override code into the scope of the main page.
@@
The injected code overrides the open and send functions of XMLHttpRequest to adapt their behavior. The open function can be modified to save properties of the connection, such as the current URL or the HTTP headers, while the send function can be adapted to dynamically customize the response callback (onreadystatechange) if the URL matches the endpoint of interest. In this function, researchers interested in fetching more data (i.e., to have a large pool of posts to re-rank) can imitate the server requests to act as the users scroll through the feed. The injected script can communicate with the browser extension’s content script using the browser’s message-passing features by broadcasting the server’s raw response. Depending on the intervention logic, when the feed needs to be modified, the customized XMLHttpRequest object can interrupt the execution, waiting for the response from the content script. Once the modified feed is ready, the request can resume by replacing the server response with the updated content. The content can be manipulated on the client side or by calling an external backend. If the logic or model can be distributed with the extension, keeping the manipulation on the client may ensure more participant privacy. Moving the edit to the backend may be preferable in other cases, like when complex logic or large models, such as LLMs, are required. Still, in this case, researchers must pay special attention to data security practices and the potential implications of sharing the data with third-party services.
Finally, it is crucial to ensure that scoring and reranking introduce minimal latency to avoid disrupting the user experience. The backend performance plays a critical role when the intervention depends on server-side logic. The implementation may require careful parallelization when relying on external services (LLM APIs) or using custom hardware (e.g., GPUs) when dedicated models are needed (e.g., fine-tuned BERT models). In specific conditions, when the model can run on the client side, researchers may consider running the models directly in the browser, potentially relying on the performance of WebGL or WASM-based implementations.
@@ 5.2. Participant recruitment
The recruitment strategy depends on many aspects, such as availability constraints, specific demographic requirements (e.g., partisans, or users of specific age groups), or length of the study.
Some experiments may aim to recruit a participant sample representative of the platform’s user base, which requires adjusting the recruitment based on the platform’s user composition (Guess, 2021). For example, in 2023, Twitter was used by 22% of the adult population with a bias toward young, highly educated, and liberals  (Center, 2024), while Facebook is more balanced on these properties but has a larger active user base of women.
Depending on the requirements and the recruitment platform, it may be desirable to first screen participants through a short survey. This preliminary screening may help exclude individuals who do not meet the eligibility criteria that the recruitment platform does not allow to filter, such as how frequently they use social media. This step is crucial to selecting the population of interest and ensuring that the study includes the right participants to receive the treatment. Researchers investigating interventions on extreme or niche content need to ensure that the recruited participants are typically exposed to the posts of interest. For example, a study found that during the 2016 election, 80% of the fake news on Twitter was consumed by only 1% of the users (Grinberg et al., 2019). If this step is necessary, we recommend keeping the survey minimal to reduce the cost and compensate participants for their time, even if they do not qualify, in accordance with common practices. Alternative methods include using ads directly on the platform of interest (Zhang et al., 2020; Rosenzweig et al., 2020), direct messages (Guess, 2021), or relying on snowball sampling with other reward incentives, such as gift cards or prizes.
@@ 5.3. Registration flow
Once the participants are recruited, the first challenge is ensuring a low-effort and scalable onboarding process. A confusing onboarding procedure can lead to high participant dropout and increase the risk of researchers being overloaded with requests for assistance.
@@
Since browser extension stores like the Chrome Store do not support direct parameter passing during extension installation (e.g., the participant ID), this section presents a solution to address this issue. This approach requires a coordination service, which can be integrated with the extension backend. It allows tracking of the user’s registration to assign the user-specific configuration to the extension that can be sourced from the recruitment platform or an initial survey.
@@
Fig. 2 summarises the registration flow using this pattern. With this flow, the recruitment platform sends the user to the coordination service directly or after an intermediary survey, which could be a pre-survey used to collect additional information necessary to run the extension (e.g., political ideology, contact information, timezone). Major survey platforms like Qualtrics support transferring information after the survey completion by combining embedded fields and redirecting URLs. The role of the coordination service is to set a persistent entry, such as an HTTP cookie, with the information that the extension must access after the installation. This page precedes the installation of the browser extension and may be used to summarize the instructions before redirecting the participant to the extension store page. Finally, when the extension is installed, opening a new tab to a page on the coordination server gives access to the HTTP cookie data previously saved that can be used to set up the extension for the current participant. This flow is simple and effective, but it is not the only approach that researchers can take. Alternative designs include integrating a custom login or implementing an OAuth flow.
Participant consent. Running an experiment that augments users’ feeds requires participant consent. Depending on the registration flow—with or without the initial survey—the participants should be presented with a description approved by an Institutional Review Board (IRB) in the case of a US university or an equivalent record in other cases. It is important to show the consent form as early as possible and interrupt the installation if the participants do not accept the terms. Some participants may perceive installing a browser extension as too invasive. In addition to the consent, we recommend including in the instruction page a clear recap of what data is and what data is not collected (e.g., “we do not have access to your direct messages”).
Error recovery. Anything (and everything) could go wrong. Especially when the goal is to run a large-scale study, it is important to implement recovery points to minimize the researchers’ workload of error fixing. For example, a participant may start the survey using a browser or a computer different from the one they plan to use during the study. It is essential to implement recovery checkpoints where the participant can recover the process. One option consists of collecting contact information such as email addresses and sending a message when they reach the instructions page. The message can contain a copy of the instructions and the link to the instructions page with all the required parameters in the URL. Alternatively, if researchers do not have access to the participants’ contact information—due to IRB restrictions or no pre-survey is planned—this step can be automated with a message on the recruitment platform.

### v2

@@ 3.3. Content Editing
In addition to reranking the content of the feed, researchers can operate directly on individual posts. This intervention can involve direct manipulation of the post, such as editing the visible social metrics (e.g., likes, comments, and shares), the appearance of the post (e.g., making some aspects of the post prominent), changing the attachments (e.g., replacing a link or a picture), or modifying the text. These interventions may be important in addressing research questions on how the platform’s design or content can affect the users. For example, the tone of existing posts can be subtly reframed to reflect a more positive sentiment or mitigate aggressive language. This intervention can allow researchers to explore how language choices in social media posts may influence user behavior, including interactions, affective experiences, discourse patterns, misinformation, or hate speech. Depending on the complexity of the desired edits, the modification can be achieved with simple dictionary substitution or complete reframing, e.g., leveraging LLMs to generate high-quality and contextual text.
Summary. This section summarises the opportunities and challenges associated with down-ranking, up-ranking, and editing feed content. Down-ranking is generally the simplest intervention, as it involves operating on content already available in the feed, whereas up-ranking, in the absence of access to the complete inventory, may require a more careful design to identify or generate posts to up-rank. Content editing can be combined with the previous methods and represents an effective strategy for testing the impact of social metrics, language, or post appearance. Furthermore, thanks to the capabilities of browser extensions, these modifications can be combined with additional interventions. Some examples include extending the platform’s functionality with widgets like the Twitter/X Community Notes labels, or adding visual indicators for certain post characteristics, such as a warning about language toxicity. These interventions can be combined to achieve more complex experimental designs that align with specific research goals.
Finally, all these interventions require researchers to pay attention to the distribution of posts in the feed of the target population, which may have repercussions for the recruitment strategy. For example, manipulating a single post may produce minimal exposure to the treatment condition, and the effect of the intervention may be too small to detect (Guess, 2021). This problem is particularly impactful in situations when users interact with only a small fraction of the content (Wang et al., 2016), typically at the top of the feed (Bakshy et al., 2015) or when the types of posts that the researchers intend to edit are rare.
@@ 4. Measurement
In this section, we describe a set of approaches for measuring the impact of the feed intervention. We summarize three alternatives that enable different insights into the effect of the feed manipulation: momentary measures, longitudinal surveys, and engagement signals.
@@ 4.1. Ecological Momentary Assessments
When researchers are interested in measuring the momentary impact of a feed intervention, a suitable design is to add surveys directly to the feed. These in-feed surveys, inspired by the approach developed in psychology called Ecological Momentary Assessment (Shiffman et al., 2008), allow researchers to capture real-time reactions directly in the context where the intervention is applied (Backstrom, 2016). In the case of social media, this strategy can elicit direct feedback on the intervention at the right time, as users scroll through their feeds. Researchers can ask highly contextualized questions such as “Are you interested in this post?” or “How does this post make you feel?”. This method can be used to measure a momentary impact of the intervention that may be too short to be measured with long-term surveys, or that is hard to measure when the question is posed without context. The insertion of these surveys can be triggered by specific events, such as when a specific post is displayed, added at regular intervals, or shown at random times while the user consumes their feed. The appearance of the survey widget may vary, from attention-grabbing designs such as popups or modal windows that can interrupt scrolling, to more subtle designs integrated as a special post in the regular feed. If integrated with the regular feed, one practical challenge of in-feed surveys is making them visible and clear enough for participants to know these questions are part of the study. If researchers require that the insertions get noticed, we recommend using colors and animations that stand out from the default design. The design must also be adapted to the different templates (e.g., dark mode) that the social media platform supports. Finally, when deciding where to place the in-feed survey, researchers must carefully consider the goals of their measurement. Participants answering the survey will be exposed to the content that precedes it in the feed, and might be influenced by other content visible in their viewport.
@@ 4.2. Survey methodology
Researchers interested in measuring the cumulative effects of the treatment can rely on a standard pre-post survey design. In such a setup, participants are presented with the same survey before (pre-) and after (post-) the experiment. Examples of attitudinal shifts that can be caused by feed interventions include changes in affective polarization, often measured using the feeling thermometer, or opinion change, typically measured with Likert-scale surveys. These surveys are assigned at the start and the end of a study and require a seamless integration into the experimental design. Pre-surveys intended to be completed at the beginning of the experiment can also be used to assign participants to specific experimental conditions, to determine whether they qualify for enrollment, or to assess systematic differences between participants who continue with the study and those who drop out. Attention checks and validation questions can be valuable tools for identifying early participants who may be less attentive and contribute unreliable data.
A similar approach, inspired by the diary study technique (Bolger et al., 2003), consists of regular assessments throughout the experiment. This method relies on administering surveys at regular intervals, e.g., daily or weekly. This design may be recommended for several reasons, such as measuring the progressive change in outcomes of interest or for a study design that requires participants to be exposed to different experimental conditions (e.g., a crossover or stepped-wedge design).
Implementing a strategy to administer surveys at regular intervals can present additional challenges, especially when the users are not online on the social media platform. Researchers may need to establish a notification system via a browser extension, implement regular messaging on the recruitment platform, or distribute a survey link through additional communication channels such as email or text messaging. This last approach requires participants’ consent to collect potentially identifiable information, which we discuss in more detail in Section 5.4. Another challenge of this setup is ensuring participants consume enough content between surveys to meaningfully measure the intervention’s impact. One potential approach, valid for repeated assessments, is to invite participants to use social media and to deliver the survey only after they have spent enough time on the platform or have been exposed to a predefined number of posts.
@@ 4.3. Engagement signals
Implementing a feed experiment using a browser extension enables researchers to go beyond self-reported measures and obtain very granular behavioral data that can provide further insights into the intervention’s effects. This includes metrics such as time spent, engagement (e.g., clicks, likes/favorites, reposts), navigation patterns, conversion rates, and other metrics (Cunningham et al., 2024) that are imperative to the business objectives of the platforms. Access to such measures also allows researchers to evaluate the trade-offs introduced by the proposed intervention. For instance, removing some content entirely (e.g., all political posts) may affect an outcome of interest (e.g., affective polarization) while also limiting disruption to engagement metrics, making adoption more attractive to social media companies.
@@ 5. Implementing a Browser Extension
In this section, we focus on the implementation details of running social media feed experiments with a browser extension and summarize various technical challenges. We offer practical recommendations and release source code  [^1] https://github.com/StanfordHCI/FeedMonitor that can be used as a blueprint to implement feed experiments on social media. This section will focus on X as a use case, but the same technological approach can be applied to any platform. Our examples are based on Google Chrome but can easily be adapted to other browsers. The rest of this section assumes familiarity with key concepts of web architecture and extension development.
@@ 5.1. Feed interventions
In this section, we present key details on implementing the browser extension’s core functionality. Although these extensions can modify many aspects of the website experience, here we specifically focus on changes related to social media feeds. Browser extensions are typically written in languages supported by the browser, such as JavaScript, HTML, and CSS, and can use various methods to modify the website experience. A common strategy involves modifying the rendered page by manipulating its Document Object Model (DOM). This method has the advantage of simplifying the identification of the element involved in the desired modification, and it gives the browser extension access to the content after the frontend scripts have already processed it. However, despite this approach being the most straightforward, for some scenarios, such as augmenting the feed, intervening at the network level may be more effective.
@@
In this case, we recommend adapting XMLHttpRequest to support the customization of the server requests. This technique offers the advantage of intercepting and modifying the incoming data before it is rendered on the page, minimizing the need for direct interaction with the page interface. Accessing the server’s communication allows logging many client-side actions shared with the server (e.g., likes and shares) without adding a listener for each event. The downside is the need to gain a deeper understanding of the communication protocol between the browser and the server. Finally, alternative approaches, such as manipulating the in-memory data structure used to render the content, are possible, but depending on the front-end framework (e.g., React), they may introduce implementation-specific complications.
@@
Implementation details. Pragmatically, to customize the behavior of the XMLHttpRequest native object, we need to run the overriding code as soon as the main page is loaded. Since behavior customization must happen within the scope of the social media page and not in the content script, we need to inject the modification script directly into the main page. This requirement necessitates a slight departure from the common pattern for developing browser extensions, as the script must be loaded in the Manifest file as a resource rather than as a content script. Then the content script can inject the override code into the main page’s scope.
@@
The injected code overrides the open and send functions of XMLHttpRequest to adapt their behavior. The open function can be modified to save properties of the connection, such as the current URL or the HTTP headers, while the send function can be adapted to dynamically customize the response callback (onreadystatechange) if the URL matches the endpoint of interest. In this function, researchers interested in fetching more data (i.e., to build a large pool of posts to rerank) can mimic server requests to simulate users scrolling through the feed. The injected script can communicate with the browser extension’s content script using the browser’s message-passing features by broadcasting the server’s raw response. Depending on the intervention logic, when the feed needs to be modified, the customized XMLHttpRequest object can pause execution while waiting for a response from the content script. Once the modified feed is ready, the request can resume by replacing the server response with the updated content. The content can be manipulated on the client side or by calling an external backend. If the logic or model can be distributed with the extension, keeping the manipulation on the client may ensure more participant privacy. Moving the edit to the backend may be preferable in other cases, like when complex logic or large models, such as LLMs, are required. Still, in this case, researchers must pay special attention to data security practices and the potential implications of sharing the data with third-party services.
Latency considerations. It is essential to ensure that scoring and reranking processes introduce minimal latency to avoid disruption of the user experience. Social media feeds are typically loaded from the server in discrete batches: for example, the X ”For You” feed loads approximately 35 items per request. Latency therefore primarily affects the moment when a user lands on the website and the feed is empty. After the initial batch of posts is rendered, platforms generally use infinite scrolling and preload subsequent batches in the background as users consume the feed. As a result, any additional latency following the initial load is largely absorbed by the platform’s existing loading mechanisms and is rarely perceptible during normal browsing.
This design substantially mitigates latency concerns because the intervention typically introduces a delay only once per session. In practice, latency requirements are therefore reduced to ensuring that the initial feed load remains within an acceptable one-time waiting threshold. Additionally, by operating within the platform’s native loading pipeline, this approach enables inheriting the platform’s strategies for managing network delays, including default loading indicators to preserve fidelity to the original user experience.
Previous research has identified several thresholds for tolerable waiting times. Studies indicate that delays of approximately 10 seconds or more often cause users to disengage or switch tasks (Nielsen, 1993). Web-focused research suggests that waiting times up to 2 seconds are generally well tolerated (Nah, 2004), with noticeable declines in user attitudes occurring beyond 4 seconds (Galletta et al., 2004). Other studies have reported upper tolerable limits of up to 12 seconds (Hoxmeier et al., 2000), depending on context and task characteristics.
Consistent with these findings, in our large-scale field study on X involving 1,256 participants (Piccardi et al., 2025), we introduced an average one-time feed-loading delay of approximately 3 seconds. In a post-study survey assessing user experience, only 24 participants (1.9%) mentioned noticing the additional latency, suggesting that delays of this magnitude are unlikely to significantly disrupt the browsing experience when they occur only during initial feed load.
When an intervention relies on server-side logic, latency arises from both network transfer and content processing. Network delay can be reduced by applying best practices in global server load balancing (GSLB). However, this optimization is often unnecessary if the majority of study participants are located within a single geographic region. In cases where the intervention can operate on the client side, network request time may be minimized by executing scripts or models directly in the browser, utilizing technologies such as WebGL or WebAssembly-based implementations. Conversely, backend performance becomes critical when the intervention is dependent on server-side logic. To maintain acceptable waiting times, implementations may require careful parallelization, particularly when using external services such as large language model APIs or using dedicated hardware, such as GPUs, for specialized models like fine-tuned BERT.
Finally, when designing an experiment, it is important to ensure that any latency is applied uniformly across experimental conditions. In particular, even in control conditions where the feed remains unchanged, the waiting time should be comparable. This can be achieved by executing the same intervention pipeline and discarding the modified output before returning it.
@@ 5.2. Participant recruitment
The recruitment strategy depends on many aspects, such as availability constraints, specific demographic requirements (e.g., partisans or users of specific age groups), or the length of the study.
Some experiments may aim to recruit a participant sample representative of the platform’s user base, which requires adjusting the recruitment based on the platform’s user composition (Guess, 2021). For example, in 2023, Twitter was used by 22% of the adult population, with a bias toward young, highly educated, and liberal users  (Center, 2024), while Facebook is more balanced on these properties but has a larger active user base of women.
Depending on the requirements and the recruitment platform, it may be desirable to first screen participants through a short survey. This preliminary screening may help exclude individuals who do not meet eligibility criteria that the recruitment platform cannot filter for, such as how frequently they use social media. This step is crucial to selecting the population of interest and ensuring that the study includes the desired participants to receive the treatment. Researchers investigating interventions on extreme or niche content need to ensure that the recruited participants are typically exposed to the posts of interest. For example, a study found that during the 2016 election, 80% of the fake news on Twitter was consumed by only 1% of the users (Grinberg et al., 2019). If this step is necessary, we recommend keeping the survey short to reduce costs and compensate participants for their time, even if they do not qualify, in accordance with common practices. Alternative methods include using ads directly on the platform of interest (Zhang et al., 2020; Rosenzweig et al., 2020), direct messages (Guess, 2021), or snowball sampling with other rewards, such as gift cards or prizes.
@@ 5.3. Registration flow
Once participants are recruited, the first challenge is to ensure a low-effort, scalable onboarding process. A confusing onboarding procedure can lead to high participant dropout and increase the risk of researchers being overloaded with requests for assistance.
@@
Since browser extension stores like the Chrome Web Store do not support passing parameters during installation (e.g., participant ID), this section presents a solution to address this issue. This approach requires a coordination service, which can be integrated with the extension backend. It allows tracking the user’s registration to assign user-specific configuration to the extension, sourced from the recruitment platform or an initial survey.
@@
Fig. 2 summarises the registration flow using this pattern. With this flow, the recruitment platform sends the user directly to the coordination service or via an intermediary survey, which could be a pre-survey used to collect additional information necessary to run the extension (e.g., political ideology, contact information, timezone). Major survey platforms like Qualtrics support transferring information after the survey completion by combining embedded fields and redirecting URLs. The role of the coordination service is to set a persistent entry, such as an HTTP cookie, containing the information the extension must access after installation. This page precedes the installation of the browser extension and may be used to summarize the instructions before redirecting the participant to the extension store page. Finally, after the extension is installed, opening a new tab to a page on the coordination server provides access to the previously saved HTTP cookie data, which can be used to set up the extension for the current participant. This flow is simple and effective, but it is not the only approach that researchers can take. Alternative designs include integrating a custom login or implementing an OAuth flow.
Participant consent. Running an experiment that augments users’ feeds requires participant consent. Depending on the registration flow—with or without the initial survey—the participants should be presented with a description approved by an Institutional Review Board (IRB) or an equivalent record in other cases. It is important to show the consent form as early as possible and interrupt the installation if the participants do not accept the terms. Some participants may perceive installing a browser extension as too invasive. In addition to the consent, we recommend including in the instruction page a clear summary of what data is and what data is not collected (e.g., “we do not have access to your direct messages”).
Error recovery. Anything (and everything) could go wrong. Especially when the goal is to run a large-scale study, it is important to implement recovery points to minimize the researchers’ workload of fixing errors. For example, a participant may start the survey using a browser or a computer different from the one they plan to use during the study. It is essential to implement recovery checkpoints so participants can recover the process. One option is to collect contact information, such as email addresses, and send a message when they reach the instructions page. The message can include a copy of the instructions and a link to the instructions page, with all required parameters in the URL. Alternatively, if researchers do not have access to participants’ contact information—due to IRB restrictions or because no pre-survey is planned—this step can be automated via a message on the recruitment platform.